OpenStudy (anonymous):

Let X be the ratio of sales to profits of some company. Assume that X has the distribution function F(x) = 0 if x < 2, F(x) = (x^2-4)/5 if 2<= x < 3, F(x) = 1 if x >= 3. Find and sketch the density. What is the probability that X is between 2.5 (40% profit) and 5 (20% profit)?

OpenStudy (anonymous):

\[F(x)=\begin{cases}0&\text{for }x<2\\\\\frac{1}{5}(x^2-4)&\text{for }2\le x<3\\\\1&\text{for }x\ge3\end{cases}\] I'll let you work on the details of the sketch. The second part asks for \(P(2.5<X<5)\), which is given by \(F(5)-F(2.5)\). Judging by the given \(F(x)\), you have \(F(5)=1\) and \(F(2.5)=\dfrac{1}{5}(2.5^2-4)=.45\), and so \(P(2.5<x<5)=.55\).